40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 72 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  13471   Wed Dec 13 09:49:23 2017 johannesUpdateASSwiring diagram

I attached a wiring schematic from the slow DAQ to the eurocrate modules. Of these, pins 1-32 (or 1A-16C) and pins 33-64 (17A-32C) are on separate DSub connectors. Therefore the easiest solution is to splice the slow DIO channels into the existing breakouts so we can proceed with the transition. This will still remove a lot of the current cable salad. For the YEND we can start thinking about a more elegant solution (For example a connector on the front panel of the Acromag chassis for the fast DIO) now that the problem is better defined.

Attachment 1: 1Y9.pdf
1Y9.pdf
  13470   Fri Dec 8 23:31:31 2017 johannesFrogsASSc1ass slow channel offloading scripts with small

While staring at epics records all day I noticed something about the PIT/YAW offset sliders and ASS offset offloading to slow channels scripts that I'm not sure others are aware off, so I'll briefly discuss it in this post.

The PIT and YAW sliders directly control soft channels that are hosted on the slow machine. Secondary epics records disentangle them for the individual coils:

  • UL = PIT+YAW
  • LL = -PIT+YAW
  • UR = PIT-YAW
  • LR = -PIT-YAW

These channels are the direct input for the physical output channels that generate the control voltage.

The fast channels for PIT and YAW have a numerical correction factor built in that accounts for differences between the OSEMs, but the slow channels don't. This means that the slow PIT/YAW controls are not entirely orthogonal but have crosstalk on the order of 10 percent. This in itself is not that dramatic, however the offload offsets scripts for the dither alignment use the fast PIT/YAW values as inputs, which represent the necessary adjustments to the OSEMs only after the individual correction factors have been applied. The offloading to slow knows nothing of this calibration difference between the OSEMs. The result is that there is a ~10 percent of the offset correction error on the mirror alignment AFTER offloading. This will of course converge after a few iterations, but in any case it is recommendable to run the dither alignment again after offloading and not offload the new offsets to the fast channels.

  13469   Fri Dec 8 12:06:59 2017 johannesOmnistructureComputersc1auxex2 ready - but need more cables

The new slow machine c1auxex2 is ready to deploy. Unfortunately we don't have enough 37pin DSub cables to connect all channels. In fact, we need a total of 8, and I found only three male-male cables and one gender changer. I asked Steve to buy more.

Over the past week I have transferred all EPICS records - soft channels and physical ones - from c1auxex to c1auxex2, making changes where needed. Today I started the in-situ testing

  1. Unplugged ETMX's satellite box
  2. Unplugged the eurocrate backplane DIN cables from the SOS Driver and QPD Whitening filter modules (the ones that receive ao channels)
  3. Measured output voltages on the relevant pins for comparison after the swap
  4. Turned off c1auxex by key, removed ethernet cable
  5. Started the modbus ioc on c1auxex2
  6. Slow machine indicator channels came online, ETMX Watchdog was responsive (but didn't have anything to do due to missing inputs) and reporting. PIT/YAW sliders function as expected
  7. Restoring the previous settings gives output voltages close to the previous values, in fact the exact values requested (due to fresh calibration)
  8. Last step is to go live with c1auxex2 and confirm the remaining channels work as expected.

I copied the relevant files to start the modbus server to /cvs/cds/caltech/target/c1auxex2, although kept local copies in /home/controls/modbusIOC/ from which they're still run.

I wonder what's the best practice for this. Probably to store the database files centrally and load them over the network on server start?

  13468   Thu Dec 7 22:24:04 2017 johannesOmnistructureComputersAcromag XEND progress

 

Quote:
 
  • Need to calibrate the +/- 10V swing of the analog channels via the USB utility, but that requires wiring the channels to the connectors and should probably be done once the unit sits in the rack
  • Need to wire power from the Sorensens into the chassis. There are +/- 5V, +/- 15V and +/- 20V present. The Acromags need only +12V-32V, for which I plan to use the +20V, and an excitation voltage for the binary channels, for which I'm going to wire the +5V. Should do this through the fuse rails on the side.
  • The current slow binary channels are sinking outputs, same as the XT1111 16-channel module we have. The additional 4 binary outputs of the XT1541 are sourcing, and I'm currently not sure if we can use them with the sos driver and whitening vme boards that get their binary control signals from the slow system.
  • Confirm switching of binary channels (haven't used model XT1111 before, but I assume the definitions are identical to XT1121)
  • Setup remaining essential EPICS channels and confirm that dimensions are the same (as in both give the same voltage for the same requested value)
  • Disconnect DIN cables, attach adapter boards + DSUB cables
  • Testing

Getting the chassis ready took a little longer than anticipated, mostly because I had not looked into the channel list myself before and forgot about Lydia's post which mentions that some of the switching controls have to be moved from the fast to the slow DAQ. We would need a total of 5+5+4+8=22 binary outputs. With the existing Acromag units we have 16 sinking outputs and 8 sourcing outputs. I looked through all the Eurocrate modules and confirmed that they all use the same switch topology which has sourcing inputs.

While one can use a pull-down resistor to control a sourcing input with a sourcing output,

pulling down the MAX333A input (datasheet says logic low is <0.8V) requires something like 100 Ohms for the pull down resistor, which would require ~150mA of current PER CHANNEL, which is unreasonable. Instead, I asked Steve to buy a second XT1111 and modified the chassis to accomodate more Acromag units.

I have now finished wiring the chassis (except for 8 remaining bypass controls to the whitening board which need the second XT1111), calibrated all channels in use, confirmed all pin locations via the existing breakout boards and DCC drawings for the eurocrate modules, and today Steve and I added more fuses to the DIN rail power distribution for +20V and +15V.

There was not enough contingent free space in the XEND rack to mount the chassis, so for now I placed it next to it.

c1auxex2 is currently hosting all original physical c1auxex channels (not yet calc records) under their original name with an _XT added at the end to avoid duplicate channel names. c1auxex is still in control of ETMX. All EPICS channels hosted by c1auxex2 are in dimensions of Volts. The plan for tomorrow is to take c1auxex off the grid, rename the c1auxex2 hosted channels and transfer ETMX controls to it, provided we can find enough 37pin DSub cables (8). I made 5 adapter boards for the 5 Eurocrate modules that need to talk to the slow DAQ through their backplane connector.

  13467   Thu Dec 7 16:28:06 2017 KojiUpdateIOOLots of red on the FE status screen

Once the RT machines were back, we launched only the five IOPs. They had bunch of red lights, but we continued to run essential models for the IFO. SOme of the lights were fixed by "global diag reset" and "mxstream restart".

The suspension were damped. We could restore the IMC lock. The locking became OK and the IMC was aligned. The REFL spot came back.

At least, I could confirm that the WFS ASC signals were not transmitted to c1mcs. There must be some disconneted links of IPC.

  13466   Thu Dec 7 15:46:31 2017 johannesHowToComputer Scripts / ProgramsLots of red on the FE status screen

[Koji, Johannes]

The issue was partially fixed and the interferometer is in workable condition now.

What -probably- fixed it was restarting the dhcp server on chiara

sudo service isc-dhcp-server restart

Afterwards the frontends were restarted one by one. SSH access was possible and the essential models for IFO operation were started.

c1iscex reported initially that no DAQ card was found, and inside the IO chassis the LED indicator strip was red. Turning off the machine, checking the cables and rebooting fixed this.

Attachment 1: 04.png
04.png
  13465   Thu Dec 7 15:02:37 2017 KojiHowToComputer Scripts / ProgramsLots of red on the FE status screen

Once a realtime machine was rebooted, it did not come back. I suspect that the diskless hosts have a difficulty to boot up.

Attachment 1: DSC_0552.JPG
DSC_0552.JPG
  13464   Thu Dec 7 11:14:37 2017 johannesHowToComputer Scripts / ProgramsLots of red on the FE status screen

Since we're getting ready to put the replacement slow DAQ for c1auxex in I wanted to bring the IFO back to operating condition after the PMC hasn't been locked for days. Something seems wrong with the CDS system though, many of the frontent models have red background and don't seem to be responsive. I followed the instructions laid out in https://wiki-40m.ligo.caltech.edu/Computer_Restart_Procedures.

In the attached screenshot, initially all c1ioo models were red, and on c1iscex only c1x01 was blue, the other ones red. I was able to ssh into both machines and tried to restart indivitual models, which didn't work and instead turned their background white. Still following the wiki page, I restarted both machines but they don't respond to pinging anymore and thus I cannot use ssh to reach them. Not sure what to do, I also rebooted fb over telnet.

So far I couldn't find any records of how to fix this situation.

Attachment 1: 22.png
22.png
  13463   Mon Dec 4 22:06:07 2017 johannesOmnistructureComputersAcromag XEND progress

I wired up the power distribution, and ethernet cables in the Acromag chassis today. For the time being it's all kind of loose in there but tomorrow the last parts should arrive from McMaster to put everything in its place. I had to unplug some of the wiring that Aaron had already done but labeled everything before I did so. I finalized the IP configuration via USB for all the units, which are now powered through the chassis and active on the network.

I started transcribing the database file ETMXaux.db that is loaded by c1auxex in the format required by the Acromags and made sure that the new c1auxex2 properly functions as a server, which it does.

ToDo-list:

  • Need to calibrate the +/- 10V swing of the analog channels via the USB utility, but that requires wiring the channels to the connectors and should probably be done once the unit sits in the rack
  • Need to wire power from the Sorensens into the chassis. There are +/- 5V, +/- 15V and +/- 20V present. The Acromags need only +12V-32V, for which I plan to use the +20V, and an excitation voltage for the binary channels, for which I'm going to wire the +5V. Should do this through the fuse rails on the side.
  • The current slow binary channels are sinking outputs, same as the XT1111 16-channel module we have. The additional 4 binary outputs of the XT1541 are sourcing, and I'm currently not sure if we can use them with the sos driver and whitening vme boards that get their binary control signals from the slow system.
  • Confirm switching of binary channels (haven't used model XT1111 before, but I assume the definitions are identical to XT1121)
  • Setup remaining essential EPICS channels and confirm that dimensions are the same (as in both give the same voltage for the same requested value)
  • Disconnect DIN cables, attach adapter boards + DSUB cables
  • Testing

 

Quote:

[Aaron, Johannes]

We configured the AtomServer for the Martian network today. Hostname is c1auxex2, IP is 192.168.113.49. Remote access over SSH is enabled.

There will be 6 acromag units served by c1auxex2.

Hostname Type IP Address
c1auxex-xt1221a 1221 192.168.113.130
c1auxex-xt1221b 1221 192.168.113.131
c1auxex-xt1221c 1221 192.168.113.132
c1auxex-xt1541a 1541 192.168.113.133
c1auxex-xt1541b 1541 192.168.113.134
c1auxex-xt1111a 1111 192.168.113.135

Some hardware to assemble the Acromag box and adapter PCBs are still missing, and the wiring and channel definitions have to be finalized. The port driver initialization instructions and channel definitions are currently locally stored in /home/controls/modbusIOC/ but will eventually be migrated to a shared location, but we need to decide how exactly we want to set up this infrastructure.

  • Should the new machines have the same hostnames as the ones they're replacing? For the transition we simply named it c1auxex2.
  • Because the communication of the server machine with the DAQ modules is happening over TCP/IP and not some VME backplane bus we could consolidate machines, particularly in the vertex area.
  • It would be good to use the fact that these SuperMicro servers have 2+ ethernet ports to separate CDS EPICS traffic from the modbus traffic. That would also keep the 30+ IPs for the Acromag thingies off the Martian host tables.
  13462   Sun Dec 3 17:01:08 2017 KojiConfigurationComputerssendmail installed on nodus

An email has come at 5PM on Dec 3rd.

 

  13461   Sun Dec 3 05:25:59 2017 gautamConfigurationComputerssendmail installed on nodus

Pizza mail didn't go out last weekend - looking at logfile, it seems like the "sendmail" service was missing. I installed sendmail following the instructions here: https://tecadmin.net/install-sendmail-server-on-centos-rhel-server/

Except that to start the sendmail service, I used systemctl and not init.d. i.e. I ran systemctl start sendmail.service (as root). Test email to myself works. Let's see if it works this weekend. Of course this isn't so critical, more important are the maintenance emails that may need to go out (e.g. disk usage alert on chiara / N2 pressure check, which looks like nodus' responsibilities). 

  13460   Fri Dec 1 17:09:29 2017 Udit KhandelwalSummaryGeneral 

Current objectives and statuses:

  • CAD Model of 40m lab (facility, chambers, invacuum components etc)
    • Status: On hold since I'm unable to acquire general dimensions of 
  13459   Thu Nov 30 10:38:39 2017 SteveUpdateVACannuloses are not pumped

Annuloses are not pumped for 30 days, since TP2 failed.  IFO pressure 7e-6 Torr it,  Rga 2.6e-6 Torr

Valve configuration: Vacuum Normal as TP3 is the forepump of Maglev, annuloses are not puped at 1.1 Torr

TP3 50K rpm,  0.15A  24C, foreline pressure 16.1 mTorr

 

Quote:

The TP3 foreline pressure was 4.8 Torr, 50K rpm 0.54A and 31C........Maglev rotation normal 560 Hz.......    IFO pressure 7.2e- 6 Torrit was not effected

V1 closed ......replaced drypump.........V1 opened

IFO 6.9e-6 Torrit at 19:55, TP3fl 18 mT,  50Krpm 0.15A 24C

VM1 is still closed

 

 

Attachment 1: AnnulosesNotPumped.png
AnnulosesNotPumped.png
  13458   Wed Nov 29 21:40:30 2017 johannesOmnistructureComputersSlow DAQ replacement computer progress

[Aaron, Johannes]

We configured the AtomServer for the Martian network today. Hostname is c1auxex2, IP is 192.168.113.49. Remote access over SSH is enabled.

There will be 6 acromag units served by c1auxex2.

Hostname Type IP Address
c1auxex-xt1221a 1221 192.168.113.130
c1auxex-xt1221b 1221 192.168.113.131
c1auxex-xt1221c 1221 192.168.113.132
c1auxex-xt1541a 1541 192.168.113.133
c1auxex-xt1541b 1541 192.168.113.134
c1auxex-xt1111a 1111 192.168.113.135

Some hardware to assemble the Acromag box and adapter PCBs are still missing, and the wiring and channel definitions have to be finalized. The port driver initialization instructions and channel definitions are currently locally stored in /home/controls/modbusIOC/ but will eventually be migrated to a shared location, but we need to decide how exactly we want to set up this infrastructure.

  • Should the new machines have the same hostnames as the ones they're replacing? For the transition we simply named it c1auxex2.
  • Because the communication of the server machine with the DAQ modules is happening over TCP/IP and not some VME backplane bus we could consolidate machines, particularly in the vertex area.
  • It would be good to use the fact that these SuperMicro servers have 2+ ethernet ports to separate CDS EPICS traffic from the modbus traffic. That would also keep the 30+ IPs for the Acromag thingies off the Martian host tables.
  13457   Wed Nov 29 15:33:16 2017 ranaUpdatePSLPMC locking

PMC wasn't locking. Had to power down c1psl. Did burt restore. Still not great.

I think many of the readbacks on the PMC MEDM screen are now bogus and misleading since the PMC RF upgrade that Gautam did awhile ago. We ought to fix the screen and clearly label which readbacks and actuators are no longer valid.

Also, the locking procedure is not so nice. The output V adjust doesn't work anymore with BLANK enabled. Would be good to make an autolocker script if we find a visitor wanting to do something fun.

  13456   Tue Nov 28 17:27:57 2017 awadeBureaucracyCalibration-RepairSR560 return, still not charging

I brought a bunch of SR560s over for repair from Bridge labs. This unit, picture attached (SN 49698), appears to still not be retaining charge. I’ve brought it back. 

Attachment 1: 96B6ABE6-CC5C-4636-902A-2E5DF553653D.jpeg
96B6ABE6-CC5C-4636-902A-2E5DF553653D.jpeg
Attachment 2: image.jpg
image.jpg
  13455   Tue Nov 28 16:02:32 2017 ranaUpdatePEMseismometer can testing

I've ordered 4 of these from McMaster. Should be delivered to the 40m by noon tomorrow.

Quote:

For the insulation, I have decided to use this one (Buna-N/PVC Foam Insulation Sheets). We will need 3 of the 1 inch plain backing ones (9349K4) to wrap a few layers around it. I'll try two layers for now, since the insulation seems to be doing quite well according to initial testing.

Kira and I also discussed the issiue. It would be good if someone can hunt aroun on the web and get some free samples of non-shedding foam with R~4.

  13454   Sun Nov 26 19:38:40 2017 SteveUpdateVACTP3 drypump replaced again

The TP3 foreline pressure was 4.8 Torr, 50K rpm 0.54A and 31C........Maglev rotation normal 560 Hz.......    IFO pressure 7.2e- 6 Torrit was not effected

V1 closed ......replaced drypump.........V1 opened

IFO 6.9e-6 Torrit at 19:55, TP3fl 18 mT,  50Krpm 0.15A 24C

VM1 is still closed

 

Attachment 1: after_replacement.png
after_replacement.png
  13453   Thu Nov 23 18:03:52 2017 gautamUpdateOptical LeversOplev "noise budget"

Here are a couple of preliminary plots of the noise from a 20minute stretch of data - the new curve is the orange one, labelled sensing, which is the spectrum of the PIT/YAW error signal from the HeNe beam single bounce off a single steering mirror onto the QPD, normalized to account for the difference in QPD sum. The peaky features that were absent in the dark noise are present here.

I am a bit confused about the total sum though - there is ~2.5mW of light incident on the PD, and the transimpedance gain is 10.7kohm. So I would expect 2.5e-3 mW * 0.4A/W * 10.7 kV/A ~ 10.7V over 4 quadrants. The ADC is 16 bit and has a range +/- 10V, so 10.7 V should be ~35,000 cts. But the observed QPD sum is ~14,000 counts. The reflected power was measured to be ~250uW, so ~10% of the total input power. Not sure if this is factored into the photodiode efficiency value of 0.4A/W. I guess there is some fraction of the QPD that doesn't generate any photocurrent (i.e. the grooves defining the quadrants), but is it reasonable that when the Oplev beam is well centered, ~50% of the power is not measured? I couldn't find any sneaky digital gains between the quadrant channels to the sum channel either... But in the Oplev setup, the QPD had ~250uW of power incident on it, and was reporting a sum of ~13,000 counts with a transimpedance gain of 100kohm, so at least the scaling seems to hold...

I guess we wan't to monitor this over a few days, see how stationary the noise profile is etc. I didn't look at the spectrum of the intensity noise during this time.

Quote:

I've setup a test setup on the ITMY Oplev table. Details + pics to follow, but for now, be aware that

Here are some pics of the setup: https://photos.app.goo.gl/DHMINAV7aVgayYcf1.None of the existing Oplev input/output steering optics were touched. Steve can make modifications as necessary, perhaps we can make similar mods to the SRM Oplev QPD and the BS one to run the HeNe test for a few days...

 

 

Attachment 1: ITMY_P_noise.pdf
ITMY_P_noise.pdf
Attachment 2: ITMY_Y_noise.pdf
ITMY_Y_noise.pdf
  13452   Wed Nov 22 23:56:14 2017 gautamUpdateOptical LeversOplev "noise budget"

Do not turn on BS/ITMY/SRM/PRM Oplev servos without reading this elog and correcting the needful!

I've setup a test setup on the ITMY Oplev table. Details + pics to follow, but for now, be aware that

  1. I've turned off the HeNe that is used for the SRM and ITMY Oplev.
  2. Moved one of the HeNe's Steve setup on the SP table to the ITMY Oplev table.
  3. Output power was 2.5mW, whereas normal power incident on this PD was ~250uW.
  4. So I changed all transimpedance gains on the ITMY Oplev QPD from 100k to 10k thin film - these should be changed back when we want to use this QPD for Oplev purposes again. Note that I did not change the compensation capacitors C3-C6, as with 10k transimpedance, and assuming they are 2.2nF, we get a corner frequency of 6.7kHz. The original schematic recommends 0.1uF. In hindsight, I should have changed these to 22nF to keep roughly the same corner frequency of ~700Hz.
    I've implemented this change as of ~5pm Nov 23 2017 - C3-C6 are now 22nF, so the corner frequency is 676Hz, as opposed to 723Hz before... This should also be undone when we use this QPD as an Oplev QPD again...
  5. I marked the position of the ITMY Oplev QPD with sharpie and also took pics so it should be easy enough to restore when we are done with this test.
  6. I couldn't get the HeNe to turn on with any of the power supplies I found in the cabinet, so I borrowed the one used to power the BS/PRM. So these oplevs are out of commission until this test is done.
  7. There is a single steering mirror in a Polaris mount which I used to center the spot on the QPD.
  8. The specular reflection (~250uW, i.e. 10% of the power incident on the QPD) is dumped onto a clean razor beam stack. Steve can put in a glass beam dump on Monday.
  9. Just in case someone accidentally turns on some servos - I've disabled the inputs to the BS, PRM and SRM oplevs, and set the limiter on the ITMY servo to 0.

Here are some pics of the setup: https://photos.app.goo.gl/DHMINAV7aVgayYcf1.None of the existing Oplev input/output steering optics were touched. Steve can make modifications as necessary, perhaps we can make similar mods to the SRM Oplev QPD and the BS one to run the HeNe test for a few days...

Quote:

too complex; just shoot straight from the HeNe to the QPD. We lower the gain of the QPD by changing the resistors; there's no sane reason to keep the existing 100k resistors for a 2 mW beam. The specular reflection of the QPD must be dumped on a black glass V dump (not some flimsy anodized aluminum or dirty razor stack)

 

  13451   Wed Nov 22 19:20:01 2017 ranaUpdateOptical LeversOplev "noise budget"

too complex; just shoot straight from the HeNe to the QPD. We lower the gain of the QPD by changing the resistors; there's no sane reason to keep the existing 100k resistors for a 2 mW beam. The specular reflection of the QPD must be dumped on a black glass V dump (not some flimsy anodized aluminum or dirty razor stack)

  13450   Wed Nov 22 17:52:25 2017 gautamUpdateOptimal ControlVisualizing cost functions

I've attempted to visualize the various components of the cost function in the way I've defined it for the current iteration of the Oplev optimal control loop design code. For each term in the cost function, the way the cost is computed depends on the ratio of the abscissa value to some threshold value (set by hand for now) - if this ratio is >1, the cost is the logarithm of the ratio, whereas if the ratio is <1, the cost is the square of the ratio. Continuity is enforced at the point at which this transition happens. I've plotted the cost function for some of the terms entering the code right now - indicated in dashed red lines are the approximate value of each of these costs for our current Oplev loop - the weights were chosen so that each of the costs were O(10) for the current controller, and the idea was that the optimizer could drive these down to hopefully O(1), but I've not yet gotten that to happen.

Based on the meeting yesterday, some possible ideas:

  1. For minimizing the control noise injection - we know the transfer function from the Oplev control signal coupling to MICH from measurements, and we also have a model for the seismic noise. So one term could be a weighted integral of (coupling - seismic) - the weight can give more importance to f>30Hz, and even more importance to f>100Hz. Right now, I don't have any suc frequency dependant requirement on the control signal.
  2. Try a simpler problem first - pendulum local damping. The position damping controller for example has fewer roots in the complex plane. Although it too has some B/R notches, which account for 16 complex roots, and hence, 32 parameters, so maybe this isn't really a simpler problem?
  3. How do we pick the number of excess poles compared to zeros in the overall transfer function? The OL loop low-pass filters are elliptic filters, which achieve the fastest transition between the passband and stopband, but for the Oplev loop roll-off, perhaps its better to have a just have some poles to roll off the HF response?
Attachment 1: globalCosts.pdf
globalCosts.pdf
  13449   Wed Nov 22 16:40:00 2017 KojiUpdateOptical LeversOplev "noise budget"

You may want to consult with the cryo Q people (Brittany, Aaron) for a Si QPD. If you want the same QPD architecture, I can look at my QPD circuit stock.

  13448   Wed Nov 22 15:29:23 2017 gautamUpdateOptical LeversOplev "noise budget"

[steve, gautam]

What is the best way to set this test up?

I think we need a QPD to monitor the spot rather than a single element PD, to answer this question about the sensor noise. Ideally, we want to shoot the HeNe beam straight at the QPD - but at the very least, we need a lens to size the beam down to the same size as we have for the return beam on the Oplevs. Then there is the power - Steve tells me we should expect ~2mW at the output of these HeNes. Assuming 100kohm transimpedance gain for each quadrant and Si responsivity of 0.4A/W at 632nm, this corresponds to 10V (ADC limit) for 250uW of power - so it would seem that we need to add some attenuating optics in the way.

Also, does anyone know of spare QPDs we can use for this test? We considered temporarily borrowing one of the vertex OL QPDs (mark out its current location on the optics table, and move it over to the SP table), but decided against it as the cabling arrangement would be too complicated. I'd like to use the same DAQ electronics to acquire the data from this test as that would give us the most direct estimate of the sensor noise for supposedly no motion of the spot, although by adding 3 optics between the HeNe and the QPD, we are introducing possible additional jitter couplings...

Quote:

For the OL NB, probably don't have to fudge any seismic noise, since that's a thing we want to suppress. More important is "what the noise would be if the suspended mirrors were no moving w.r.t. inertial space".

For that, we need to look at the data from the OL test setup that Steve is putting on the SP table.

 

Attachment 1: OplevTest.jpg
OplevTest.jpg
  13447   Wed Nov 22 14:47:03 2017 KiraUpdatePEMseismometer can testing

For the insulation, I have decided to use this one (Buna-N/PVC Foam Insulation Sheets). We will need 3 of the 1 inch plain backing ones (9349K4) to wrap a few layers around it. I'll try two layers for now, since the insulation seems to be doing quite well according to initial testing.

Quote:

Updated some values, most importantly, the k-factor. I had assumed that it was in the correct units already, but when converting it to 0.046 W/(m^2*K) from 0.26 BTU/(h*ft^2*F), I got the following plot. The time constant is still a bit larger than what we'd expect, but it's much better with these adjustments.

For our next steps, I will measure the time constant of the heater without any insulation and then decide how many layers of it we will need. I'll need to construct and calibrate a temperature sensor like the ones I've made before and use it to record the values more accurately.

 

 

  13446   Wed Nov 22 12:13:15 2017 KiraUpdatePEMseismometer can testing

Updated some values, most importantly, the k-factor. I had assumed that it was in the correct units already, but when converting it to 0.046 W/(m^2*K) from 0.26 BTU/(h*ft^2*F), I got the following plot. The time constant is still a bit larger than what we'd expect, but it's much better with these adjustments.

For our next steps, I will measure the time constant of the heater without any insulation and then decide how many layers of it we will need. I'll need to construct and calibrate a temperature sensor like the ones I've made before and use it to record the values more accurately.

Quote:

I performed a test with the can last week with one layer of insulation to see how well it worked. First, I soldered two heaters together in series so that the total resistance was 48.6 ohms. I placed the heaters on the sides of the can and secured them. Then I wrapped the sides and top of the can in insulation and sealed the edges with tape, only leavng the handles open. I didn't insulate the bottom. I connected the two ends of the heater directly into the DC source and drove the current as high as possible (around 0.6A). I let the can heat up to a final value of 37.5C, turned off the current and manually measured the temperature, recoding the time every half degree. I then plotted the results, along with a fit. The intersection of the red line with the data marks the time constant and the temperature at which we get the time constant. This came out to be about 1.6 hours, much longer than expected considering that onle one layer instead of four was used. With only one layer, we would expect the time constant to be about 13 min, while for 4 layers it should be 53 min (the area A is 0.74 m^2 and not 2 m^2).

Quote:

I made a model for our seismometer can using actual data so that we know approximately what the time constant should be when we test it out. I used the appendix in Megan Kelley's report to make a relation for the temperature in terms of time.

\frac{dQ}{dt}=mc\frac{dT(t)}{dt} so T(t)=\frac{1}{mc}\int \frac{dQ}{dt}dt=\frac{1}{mc}\int P_{net}dt and P_{net}=P_{in}-P_{out}

In our case, we will heat the can to a certain temerature and wait for it to cool on its own so P_{in}=0

We know that P_{out}=\frac{kA\Delta T}{d} where k is the k-factor of the insulation we are using, A is the area of the surface through which heat is flowing, \Delta T is the change in temperature, d is the thickness of the insulation.

Therefore,

T(t)=\frac{1}{mc}\int_{0}^{t}\frac{kA}{d}[T_{lab}-T(t')]dt'=\frac{kA}{mcd}(T_{lab}t-\int_{0}^{t}T(t')dt')

We can take the derivative of this to get

T'(t)=\frac{kAT_{lab}}{mcd}-\frac{kA}{mcd}T(t), or T'(t)=B-CT(t) 

We can guess the solution to be

T(t)=C_{1}e^{-t/\tau}+C_{2} where tau is the time constant, which we would like to find.

The boundary conditions are T(0)=40 and T(\infty)=T_{lab}=24. I assumed we would heat up the can to 40 celcius while the room temp is about 24. Plugging this into our equations,

C_{1}+C_{2}=40, C_{2}=24, so C_{1}=16

We can plug everything back into the derivative T'(t)

T'(t)=-\frac{16}{\tau}e^{-t/\tau}=B-C[16e^{-t/\tau}+24]

Equating the exponential terms on both sides, we can solve for tau

\frac{16}{\tau}e^{-t/\tau}=16Be^{-t/\tau}, \frac{1}{\tau}=B, \tau=\frac{1}{B}=\frac{mcd}{kA}

Plugging in the values that we have, m = 12.2 kg, c = 500 J/kg*k (stainless steel), d = 0.1 m, k = 0.26 W/(m^2*K), A = 2 m^2, we get that the time constant is 0.326hr. I have attached the plot that I made using these values. I would expect to see something similar to this when I actually do the test.

To set up the experiment, I removed the can (with Steve's help) and will place a few heating pads on the outside and wrap the whole thing in a few layers of insulation to make the total thickness 0.1m. Then, we will attach the heaters to a DC source and heat the can up to 40 celcius. We will wait for it to cool on its own and monitor the temperature to create a plot and find the experimental time constant. Later, we can use the heatng circuit we used for the PSL lab and modify the parts as needed to drive a few amps through the circuit. I calculated that we'd need about 6A to get the can to 50 celcius using the setup we used previously, but we could drive a smaller current by using a higher heater resistance.

 

 

Attachment 1: cooling_fit_1.png
cooling_fit_1.png
  13445   Wed Nov 22 11:51:38 2017 gautamConfigurationComputersnodus post OS migration admin

Confirmed that this crontab is running - the daily backup of the crontab seems to have successfully executed, and there is now a file crontab_nodus.ligo.caltech.edu.20171122080001 in the directory quoted below. The $HOSTNAME seems to be "nodus.ligo.caltech.edu" whereas it was just "nodus", so the file names are a bit longer now, but I guess that's fine...

Quote:

I restored the nodus crontab (copied over from the Nov 17 backup of the same at /opt/rtcds/caltech/c1/scripts/crontab/crontab_nodus.20171117080001. There wasn't a crontab, so I made one using sudo crontab -e.

This crontab is supposed to execute some backup scripts, send pizza emails, check chiara disk usage, and backup the crontab itself.

I've commented out the backup of nodus' /etc and /export for now, while we get back to fully operational nodus (though we also have a backup of /cvs/cds/caltech/nodus_backup on the external LaCie drive), they can be re-enabled by un-commenting the appropriate lines in the crontab.

 

 

  13444   Wed Nov 22 05:41:32 2017 ranaUpdateOptical LeversOplev "noise budget"

For the OL NB, probably don't have to fudge any seismic noise, since that's a thing we want to suppress. More important is "what the noise would be if the suspended mirrors were no moving w.r.t. inertial space".

For that, we need to look at the data from the OL test setup that Steve is putting on the SP table.

  13443   Wed Nov 22 00:54:18 2017 johannesOmnistructureComputersSlow DAQ replacement computer progress

I got the the SuperMicro 1U server box from Larry W on Monday and set it up in the CryoLab for initial testing.

The specs: https://www.supermicro.com/products/system/1U/5015/SYS-5015A-EHF-D525.cfm

The processor is an Intel D525 dual core atom processor with 1.8 GHz (i386 architecture, no 64-bit support). The unit has a 250GB SSD and 4GB RAM.

I installed Debian Jessie on it without any problems and compiled the most recent stable versions of EPICS base (3.15.5), asyn drivers (4-32), and modbus module (2-10-1). EPICS and asyn each took about 10 minutes, and modbus about 1 minute.

I copied the database files and port driver definitions for the cryolab from cryoaux, whose modbus services I suspended, and initialized the EPICS modbus IOC on the SuperMicro machine instead. It's working flawlessly so far, but admittedly the box is not under heavy load in the cryolab, as the framebuilder there is logging only the 16 analog channels.

I have recently worked out some kinks in the port driver and channel definitions, most importantly:

  • mosbus IOC initialization is performed automatically by systemd on reboot
  • If the IOC crashes or a system reboot is required the Acromag units freeze in their last current state. When the IOC is started a single read operation of all A/D registers is performed and the result taken as the initial value of the corresponding channel, causing no discontinuity in generated voltage EVER (except of course for the rare case when the Acromags themselves have to be restarted)

Aaron and I set 12/4 as a tentative date when we will be ready to attempt a swap. Until then the cabling needs to be finished and a channel database file needs to be prepared.

  13442   Tue Nov 21 23:47:51 2017 gautamConfigurationComputersnodus post OS migration admin

I restored the nodus crontab (copied over from the Nov 17 backup of the same at /opt/rtcds/caltech/c1/scripts/crontab/crontab_nodus.20171117080001. There wasn't a crontab, so I made one using sudo crontab -e.

This crontab is supposed to execute some backup scripts, send pizza emails, check chiara disk usage, and backup the crontab itself.

I've commented out the backup of nodus' /etc and /export for now, while we get back to fully operational nodus (though we also have a backup of /cvs/cds/caltech/nodus_backup on the external LaCie drive), they can be re-enabled by un-commenting the appropriate lines in the crontab.

Quote:

The post OS migration admin for nodusa bout apache, elogd, svn, iptables, etc can be found in https://wiki-40m.ligo.caltech.edu/NodusUpgradeNov2017

Update: The svn dump from the old svn was done, and it was imported to the new svn repository structure. Now the svn command line and (simple) web interface is running. "websvn" is not installed.

 

  13441   Tue Nov 21 23:04:12 2017 gautamUpdateOptical LeversOplev "noise budget"

Per our discussions in the meetings over the last week, I've tried to put together a simple Oplev noise budget. The only two terms in this for now are the dark noise and a model for the seismic noise, and are plotted together with the measured open-loop error signal spectra.

  1. Dark noise
    • Beam was taken off the OL QPD
    • A small DC offset was added to all the oplev segment input filters to make the sum ~20-30 cts [call this testSum] (usually it varies from 4000-13000 for the BS/ITMs, call this nominalSum).
    • I downloaded 20mins of dq-ed error signal data, and computed the ASD, dividing by a factor of nominalSum / testSum to account for the usual light intensity on the QPD.
  2. Seismic noise
    • This is a very simplistic 1/f^2 pendulum TF with a pair of Q=2 poles at 1Hz.
    • I adjusted the overall gain such that the 1Hz peak roughly line up in measurement and model.
    • The stack isn't modelled at all.

Some remarks:

  • The BS oplev doesn't have any whitening electronics, and so has a higher electronics noise floor compared to the ITMs. But it doesn't look like we are limited by this lower noise floor anywhere..
  • I wonder what all those high frequency features seen in the ITM error signal spectra are - mechanical resonances of steering optics? It is definitely above the dark noise floor, so I am inclined to believe this is real beam motion on the QPD, but surely this can't be the test-mass motion? If it were, the measured A2L would be much higher than the level it is adjudged to be at now. Perhaps it's some resonances of steering mirrors?
  • The seismic displacement @100Hz per the GWINC model is ~1e-19 m/rtHz. Assuming the model A2L = d_rms * theta(f) where d_rms is the rms offset of the beam spot from the optic center, and theta(f) is the angular control signal to the optic, for a 5mm rms offset of the spot from the center, theta(f) must be ~1e-17 urad @100Hz. This gives some requirement on the low pass required - I will look into adding this to the global optimization cost.

 

Attachment 1: vertexOL_noises.pdf
vertexOL_noises.pdf vertexOL_noises.pdf vertexOL_noises.pdf
  13440   Tue Nov 21 17:51:01 2017 KojiConfigurationComputersnodus post OS migration admin

The post OS migration admin for nodusa bout apache, elogd, svn, iptables, etc can be found in https://wiki-40m.ligo.caltech.edu/NodusUpgradeNov2017

Update: The svn dump from the old svn was done, and it was imported to the new svn repository structure. Now the svn command line and (simple) web interface is running. And "websvn" was also implemented.

  13439   Tue Nov 21 16:28:23 2017 gautamUpdateOptical LeversBS OL calibration updated

The numbers I have from the fitting don't agree very well with the OSEM readouts. Attachment #1 shows the Oplev pitch and yaw channels, and also the OSEM ones, while I swept the ASC_PIT offset. The output matrix is the "naive" one of (+1,+1,-1,-1). SUSPIT_IN1 reports ~30urad of motion, while SUSYAW_IN1 reports ~10urad of motion.

From the fits, the BS calibration factors were ~x8 for pitch and x12 for yaw - so according to the Oplev channels, the applied sweep was ~80urad in pitch, and ~7urad in yaw.

Seems like either (i) neither the Oplev channels nor the OSEMs are well diagonalized and that their calibration is off by a factor of ~3 or (ii) there is some significant imbalance in the actuator gains of the BS coils...

Quote:
 

Need to double check against OSEM readout during the sweep.

 

Attachment 1: BS_oplev_sweep.png
BS_oplev_sweep.png
  13438   Tue Nov 21 16:00:05 2017 KiraUpdatePEMseismometer can testing

I performed a test with the can last week with one layer of insulation to see how well it worked. First, I soldered two heaters together in series so that the total resistance was 48.6 ohms. I placed the heaters on the sides of the can and secured them. Then I wrapped the sides and top of the can in insulation and sealed the edges with tape, only leavng the handles open. I didn't insulate the bottom. I connected the two ends of the heater directly into the DC source and drove the current as high as possible (around 0.6A). I let the can heat up to a final value of 37.5C, turned off the current and manually measured the temperature, recoding the time every half degree. I then plotted the results, along with a fit. The intersection of the red line with the data marks the time constant and the temperature at which we get the time constant. This came out to be about 1.6 hours, much longer than expected considering that onle one layer instead of four was used. With only one layer, we would expect the time constant to be about 13 min, while for 4 layers it should be 53 min (the area A is 0.74 m^2 and not 2 m^2).

Quote:

I made a model for our seismometer can using actual data so that we know approximately what the time constant should be when we test it out. I used the appendix in Megan Kelley's report to make a relation for the temperature in terms of time.

\frac{dQ}{dt}=mc\frac{dT(t)}{dt} so T(t)=\frac{1}{mc}\int \frac{dQ}{dt}dt=\frac{1}{mc}\int P_{net}dt and P_{net}=P_{in}-P_{out}

In our case, we will heat the can to a certain temerature and wait for it to cool on its own so P_{in}=0

We know that P_{out}=\frac{kA\Delta T}{d} where k is the k-factor of the insulation we are using, A is the area of the surface through which heat is flowing, \Delta T is the change in temperature, d is the thickness of the insulation.

Therefore,

T(t)=\frac{1}{mc}\int_{0}^{t}\frac{kA}{d}[T_{lab}-T(t')]dt'=\frac{kA}{mcd}(T_{lab}t-\int_{0}^{t}T(t')dt')

We can take the derivative of this to get

T'(t)=\frac{kAT_{lab}}{mcd}-\frac{kA}{mcd}T(t), or T'(t)=B-CT(t) 

We can guess the solution to be

T(t)=C_{1}e^{-t/\tau}+C_{2} where tau is the time constant, which we would like to find.

The boundary conditions are T(0)=40 and T(\infty)=T_{lab}=24. I assumed we would heat up the can to 40 celcius while the room temp is about 24. Plugging this into our equations,

C_{1}+C_{2}=40, C_{2}=24, so C_{1}=16

We can plug everything back into the derivative T'(t)

T'(t)=-\frac{16}{\tau}e^{-t/\tau}=B-C[16e^{-t/\tau}+24]

Equating the exponential terms on both sides, we can solve for tau

\frac{16}{\tau}e^{-t/\tau}=16Be^{-t/\tau}, \frac{1}{\tau}=B, \tau=\frac{1}{B}=\frac{mcd}{kA}

Plugging in the values that we have, m = 12.2 kg, c = 500 J/kg*k (stainless steel), d = 0.1 m, k = 0.26 W/(m^2*K), A = 2 m^2, we get that the time constant is 0.326hr. I have attached the plot that I made using these values. I would expect to see something similar to this when I actually do the test.

To set up the experiment, I removed the can (with Steve's help) and will place a few heating pads on the outside and wrap the whole thing in a few layers of insulation to make the total thickness 0.1m. Then, we will attach the heaters to a DC source and heat the can up to 40 celcius. We will wait for it to cool on its own and monitor the temperature to create a plot and find the experimental time constant. Later, we can use the heatng circuit we used for the PSL lab and modify the parts as needed to drive a few amps through the circuit. I calculated that we'd need about 6A to get the can to 50 celcius using the setup we used previously, but we could drive a smaller current by using a higher heater resistance.

 

Attachment 1: cooling_fit.png
cooling_fit.png
Attachment 2: IMG_20171121_164835.jpg
IMG_20171121_164835.jpg
  13437   Tue Nov 21 11:37:29 2017 gautamUpdateOptical LeversBS OL calibration updated

I calibrated the BS oplev PIT and YAW error signals as follows:

  1. Locked X-arm, ran dither alignment servos to maximize transmission.
  2. Applied an offset to the ASC PIT/YAW filter banks. Set the ramp time to something long, I used 60 seconds.
  3. Monitored the X arm transmission while the offset was being ramped, and also the oplev error signal with its current calibration factor.
  4. Fit the data, oplev error signal vs arm transmission, with a gaussian, and extracted the scaling factor (i.e. the number which the current Oplev error signals have to be multiplied by for the error signal to correspond to urad of angular misalignment as per the overlap of the beam axis to the cavity axis.
  5. Fits are shown in Attachment #1 and #2.
  6. I haven't done any error analysis yet, but the open loop OL spectra for the BS now line up better with the other optics, see Attachment #3 (although their calibration factors may need to be updated as well...). Need to double check against OSEM readout during the sweep.
  7. New numbers have been SDF-ed.

The numbers are:

BS Pitch     15  /  130    (old/new)     urad/counts

BS Yaw       14  /  170    (old/new)     urad/counts

Quote:

I bet the calibration is out of date; probably we replaced the OL laser for the BS and didn't fix the cal numbers. You can use the fringe contrast of the simple Michelson to calibrate the OLs for the ITMs and BS.

 

Attachment 1: OL_calib_BS_PERROR.pdf
OL_calib_BS_PERROR.pdf
Attachment 2: OL_calib_BS_YERROR.pdf
OL_calib_BS_YERROR.pdf
Attachment 3: VertexOLnoise_updated.pdf
VertexOLnoise_updated.pdf
  13436   Tue Nov 21 11:21:26 2017 gautamUpdateCDSRFM network down

I noticed yesterday evening that I wasn't able to engage the single arm locking servos - turned out that they weren't getting triggered, which in turn pointed me to the fact that the arm transmssion channels seemed dead. Poking around a little, I found that there was a red light on the CDS overview screen for c1rfm.

  • The error seems to be in the receiving model only, i.e. c1rfm, all the sending models (e.g. c1scx) don't report any errors, at least on the CDS overview screen.
  • Judging by dataviewer trending of the c1rfm status word, seems like this happened on Sunday morning, around 11am.
  • I tried restarting both sender and receiver models, but error persists.
  • I got no useful information from the dmesg logs of either c1sus (which runs c1rfm), or c1iscex (which runs c1scx).
  • There are no physical red lights in the expansion chassis that I could see - in the past, when we have had some timing errors, this would be a signature.

Not sure how to debug further...

* Fix seems to be to restart the sender RFM models (c1scx, c1scy, c1asx, c1asy).

Attachment 1: RFMerrors.png
RFMerrors.png
  13435   Fri Nov 17 17:10:53 2017 ranaOmnistructureComputersAcromag wired up

Exactly: you'll have to list explicitly what functions those channels had so that we know what we're losing before we make the switch.

  13434   Fri Nov 17 16:31:11 2017 aaronOmnistructureComputersAcromag wired up

Acromag Wireup Update

I finished wiring up the Acromags to replace the VME boxes on the x arm. I still need to cut down the bar and get them all tidy in the box, but I wanted to post the wiring maps I made.
I wanted to note specifically that a few of the connections were assigned to VME boxes but are no longer assigned in this Acromag setup. We should be sure that we actually do not need to use the following channels:

Channels no longer in use

  • From the VME analog output (VMIVME 4116) to the QPD Whitening board (no DCC number on the front), 3 channels are no longer in use
  • From the anti-image filter (D000186) to the ADC (VMIVME 3113A) 5 channels are no longer in use (these are the only channels from the anti-image filter, so this filter is no longer in use at all?)
  • From the universal dewhitening filter (D000183) to a binary I/O adapter (channels 1-16), 4 channels are no longer in use. These are the only channels from the dewhitening filter
  • From a second universal dewhitening filter (D000183) to another the binary I/O adapter (channels 1-16), one channel is no longer in use (this was the only channel from this dewhitening filter).
  • From the opti-lever (D010033) to the VME ADC (VMIVME 3113A), 7 channels are no longer in use (this was all of the channels from the opti lever)
  • From the SUS PD Whitening/Interface board (D000210) to a binary I/O adapter (channels 1-16), 5 channels are no longer in use. 
  • Note that none of the binary I/O adapter channels are in use.

 

Attachment 1: AcromagWiringMaps.pdf
AcromagWiringMaps.pdf AcromagWiringMaps.pdf AcromagWiringMaps.pdf AcromagWiringMaps.pdf AcromagWiringMaps.pdf AcromagWiringMaps.pdf AcromagWiringMaps.pdf
  13433   Thu Nov 16 15:43:01 2017 ranaUpdateOptical LeversOptical lever noise

I bet the calibration is out of date; probably we replaced the OL laser for the BS and didn't fix the cal numbers. You can use the fringe contrast of the simple Michelson to calibrate the OLs for the ITMs and BS.

  13432   Thu Nov 16 13:57:01 2017 gautamUpdateOptical LeversOptical lever noise

I disabled the OL loops for ITMX, ITMY and BS at GPStime 1194897655 to come up with an Oplev noise budget. OL spots were reasonably well centered - by that, I mean that the PIT/YAW error signals were less than 20urad in absolute value.

Attachment #1 is a first look at the DTT spectra - I wonder why the BS Oplev signals don't agree with the ITMs at ~1Hz? Perhaps the calibration factor is off? The sensing noise not really flat above 100Hz - I wonder what all those peaky features are. Recall that the ITM OLs have analog whitening filters before the ADC, but the BS doesn't...

In Attachment #2, I show comparison of the error signal spectra for ITMY and SRM - they're on the same stack, but the SRM channels don't have analog de-whitening before the ADC.

For some reason, DTT won't let me save plots with latex in the axes labels...

Attachment 1: VertexOLnoise.pdf
VertexOLnoise.pdf
Attachment 2: ITMYvsSRM.pdf
ITMYvsSRM.pdf
  13431   Thu Nov 16 00:53:26 2017 gautamUpdateLSCDRMI noise sub-budgets

I've incorporated the functionality to generate sub-budgets for the various grouped traces in the NBs (e.g. the "A2L" trace is really the quadrature sum of the A2L coupling from 6 different angular servos).

For now, I'm only doing this for the A2L coupling, and the AUX length loop coupling groups. But I've set up the machinery in such a way that doing so for more groups is easy.

Here are the sub-budget plots for last night's lock - for the OL plot, there are only 3 lines (instead of 6) because I group the PIT and YAW contributions in the function that pulls the data from the nds server, and don't ever store these data series individually. This should be rectified, because part of the point of making these sub-budgets is to see if there is a particularly bad offender in a given group.

I'll do a quick OL loop noise budget for the ITM loops tomorrow.

I also wonder if it is necessary to measure the Oplev A2L coupling from lock to lock? This coupling will be dependant on the spot position on the optic, and though I run the dither alignment servos to minimize REFL_DC, AS_DC, I don't have any intuition for how the offset from center of optic varies from lock to lock, and if this is at all significant. I've been using a number from a measurement made in May. Need to do some algebra...

Attachment 1: C1NB_a2l_40m_MICH_NB_2017-11-15.pdf
C1NB_a2l_40m_MICH_NB_2017-11-15.pdf
Attachment 2: C1NB_aux_40m_MICH_NB_2017-11-15.pdf
C1NB_aux_40m_MICH_NB_2017-11-15.pdf
  13430   Thu Nov 16 00:45:39 2017 gautamUpdateSUSSOS Sapphire Prism design

 

Quote:
 
  • If I could get pictures of the lower mirror clamp (document D960008), it would be helpful in making solidworks model. Document is unclear. Same for sensor/actuator head assembly. 

If you go through this thread of elogs, there are lots of pictures of the SOS assembly with the optic in it from the vent last year. I think there are many different perspectives, close ups of the standoffs, and of the OSEMs in their holders in that thread.

This elog has a measurement of the pendulum resonance frequencies with ruby standoffs - although the ruby standoff used was cylindrical, and the newer generation will be in the shape of a prism. There is also a link in there to a document that tells you how to calculate the suspension resonance frequencies using analytic equations.

  13429   Thu Nov 16 00:14:47 2017 Udit KhandelwalUpdateSUSSOS Sapphire Prism design

Summary:

  • SOS solidworks model is nearly complete
    • Having trouble with the design of the sensor/actuator head assembly and the lower clamps
  • After Gautam's suggestion, installed Abaqus on computer. Teaching it to myself to eventually do FEM analysis and find resonant frequency of the system
    • Goal is to replicate frequency listed in the SOS documents to confirm accuracy of computer model, then replace guide rods with sapphire prisms and change geometry to get same results

 

Questions:

  • How accurate do the details (like fillet, chamfer, placement of little vent holes), and material of the different SOS parts need to be in the model?
  • If I could get pictures of the lower mirror clamp (document D960008), it would be helpful in making solidworks model. Document is unclear. Same for sensor/actuator head assembly. 
  13428   Wed Nov 15 01:37:07 2017 gautamUpdateLSCDRMI low freq. nosie improved

Pianosa just crashed and ate my elog, along with all the DTT/Foton windows I had open, so more details tomorrow... This workstation has been crashing ~once a month for the last 6 months.

Summary:

Below ~100Hz, the hypothesis is that the BS oplev A2L contribution dominates the MICH displacement noise. I wanted to see if I could mitigate this my modifying the BS Oplev loop shape.

Details:

  • Swept sine TF measurements suggested that the BS A2L contribution is between 10-100x that of the ITM A2L
  • The Oplev loop shape for BS is different from ITMs - specifically, there is a Res-gain centered at ~3.3 Hz. The low frequency ~0.6Hz boost filter present in the ITM Oplev loops was disengaged for the BS Oplves.
  • I turned off the BS OL loops and looked at error signal spectra - didn't seem that different from ITM OL error signals, so I decided to try turning off the res-gain and engage the 0.6Hz boost.
  • This change also gave me much more phase at ~6Hz, which is roughly the UGF of the loop. So I put in another roll-off low pass filter with corner frequency 25Hz. 
  • This worked okay - RMS went down by ~5x (which is even better than the original config), and although the performance between ~3 and 10Hz is slightly worse than with the old combination,this region isn't the dominant contribution to the RMS. PM at the upper UGF is ~30degrees in the new configuration. 
  • I wanted to give DRMI locking a shot with the new OL loop - expectations were that the noise between 30-100Hz would improve, and perhaps the engaging of de-whitening filters on BS would also be easier given the more severe roll-off at high-frequencies.
  • Attachment #1 shows the NB for tonights lock. All MICH optics had their coil drivers de-whitened, and all the LSC PDs were whitened for this measurement.
  • I've edited the NB code to make the A2L calculation more straightforward, I now just make the coupling 1/f^2 and give the function a measured overall gain, so that this curve can now be easily added to all future NBs. I've also transcribed the matlab funciton used for parsing Foton files into python, this allows me to convert the DQ-ed OL error signals to control signals. Will update git with changes.

Remarks:

  1. MICH noise has improved by ~2x between 40-80Hz.
  2. Not sure what to make of the broad hump around 60Hz - scatter shelf?
  3. There is still unexplained noise below 100Hz, the A2L estimate is considerably lower than the measured noise.
  4. We are still more than an order of magnitude away from the estimated seismic noise floor at low frequencies (but getting closer!).

I've been banging my head against optimal loop shaping, with the OL loop as a test-case, without much success - as was the case with coating PSO, the magic is in smartly defining the cost function, but right now, my optimizer seems to be pushing most of the roots I'm making available for it to place to high frequencies. I've got a term in there that is supposed to guard against this, need to tweak further...


Attachment #2: Eye-fits of measured OL A2L coupling TFs to a 1/f^2 shape, with the gain being the parameter "fitted". I used these value, and the DQ-ed OL error signal in lock, to estimate the red curve labelled "A2L" in Attachment #1. The dots are the measurement, and the lines are the 1/f^2 estimates.

Attachment 1: C1NB_disp_40m_MICH_NB_2017-11-15.pdf
C1NB_disp_40m_MICH_NB_2017-11-15.pdf
Attachment 2: OL_A2L_couplings.pdf
OL_A2L_couplings.pdf
  13427   Tue Nov 14 16:02:43 2017 KiraSummaryPEMseismometer can testing

I made a model for our seismometer can using actual data so that we know approximately what the time constant should be when we test it out. I used the appendix in Megan Kelley's report to make a relation for the temperature in terms of time.

\frac{dQ}{dt}=mc\frac{dT(t)}{dt} so T(t)=\frac{1}{mc}\int \frac{dQ}{dt}dt=\frac{1}{mc}\int P_{net}dt and P_{net}=P_{in}-P_{out}

In our case, we will heat the can to a certain temerature and wait for it to cool on its own so P_{in}=0

We know that P_{out}=\frac{kA\Delta T}{d} where k is the k-factor of the insulation we are using, A is the area of the surface through which heat is flowing, \Delta T is the change in temperature, d is the thickness of the insulation.

Therefore,

T(t)=\frac{1}{mc}\int_{0}^{t}\frac{kA}{d}[T_{lab}-T(t')]dt'=\frac{kA}{mcd}(T_{lab}t-\int_{0}^{t}T(t')dt')

We can take the derivative of this to get

T'(t)=\frac{kAT_{lab}}{mcd}-\frac{kA}{mcd}T(t), or T'(t)=B-CT(t) 

We can guess the solution to be

T(t)=C_{1}e^{-t/\tau}+C_{2} where tau is the time constant, which we would like to find.

The boundary conditions are T(0)=40 and T(\infty)=T_{lab}=24. I assumed we would heat up the can to 40 celcius while the room temp is about 24. Plugging this into our equations,

C_{1}+C_{2}=40, C_{2}=24, so C_{1}=16

We can plug everything back into the derivative T'(t)

T'(t)=-\frac{16}{\tau}e^{-t/\tau}=B-C[16e^{-t/\tau}+24]

Equating the exponential terms on both sides, we can solve for tau

\frac{16}{\tau}e^{-t/\tau}=16Be^{-t/\tau}, \frac{1}{\tau}=B, \tau=\frac{1}{B}=\frac{mcd}{kA}

Plugging in the values that we have, m = 12.2 kg, c = 500 J/kg*k (stainless steel), d = 0.1 m, k = 0.26 W/(m^2*K), A = 2 m^2, we get that the time constant is 0.326hr. I have attached the plot that I made using these values. I would expect to see something similar to this when I actually do the test.

To set up the experiment, I removed the can (with Steve's help) and will place a few heating pads on the outside and wrap the whole thing in a few layers of insulation to make the total thickness 0.1m. Then, we will attach the heaters to a DC source and heat the can up to 40 celcius. We will wait for it to cool on its own and monitor the temperature to create a plot and find the experimental time constant. Later, we can use the heatng circuit we used for the PSL lab and modify the parts as needed to drive a few amps through the circuit. I calculated that we'd need about 6A to get the can to 50 celcius using the setup we used previously, but we could drive a smaller current by using a higher heater resistance.

Attachment 1: time_const.png
time_const.png
  13426   Tue Nov 14 08:54:37 2017 SteveUpdateIOOMC1 glitching
Attachment 1: MC1_glitching.png
MC1_glitching.png
  13425   Fri Nov 10 18:57:41 2017 ranaSummaryElectronicsIthaca 1201 vs. SR560

I characterized the black Ithaca 1201 pre-amp that we had sitting in the racks. It works fine and the input referred noise is < 10 nV/rHz. I also checked that the filter selection switches on the front panel did what they claim and that the gain knob gives us the correct gain.

For comparison I have also included the G=100, 1000 input referred noise of one of the best SR560 that we have (s/n 02763) in the lab. Above a few Hz, the SR560 is better, but for low frequency measurements it seems that the 1201 is our friend.

As with the SR560, you don't actually get low noise performance for G < 100, due to some fixed output noise level.

Steve:  sn48332 of Ithaca 1201

Attachment 1: Ithaca1201.pdf
Ithaca1201.pdf
  13424   Fri Nov 10 13:46:26 2017 SteveUpdatePEMM3.1 local earthquake

BOOM SHAKA-LAKA

Attachment 1: 3.1M_local_EQ.png
3.1M_local_EQ.png
  13423   Fri Nov 10 08:52:21 2017 SteveUpdateVACTP3 drypump replaced

PSL shutter closed at 6e-6 Torr-it    

   The foreline pressure of the drypump is 850 mTorr at 8,446 hrs of seal life

V1 will be closed for ~20 minutes for drypump replacement..........

9:30am dry pump replaced, PSL shutter opened at 7.7E-6 Torr-it

  Valve configuration: vacuum normal as  TP3 is the forepump of the Maglev  & annuloses are not pumped.

Quote:

TP3 drypump replaced at 655 mTorr, no load, tp3 0.3A 

This seal lasted only for 33 days at  123,840 hrs

The replacement is performing well: TP3 foreline pressure is 55 mTorr, no load, tp3 0.15A at 15 min  [ 13.1 mTorr at d5 ]

 

Valve configuration: Vacuum Normal, ITcc 8.5E-6 Torr

Quote:

Dry pump of TP3 replaced after 9.5 months of operation.[ 45 mTorr d3 ]

The annulosses are pumped.

Valve configuration: vac normal, IFO pressure 4.5E-5 Torr [1.6E-5 Torr d3 ] on new ITcc gauge, RGA is not installed yet.

Note how fast the pressure is dropping when the vent is short.

Quote:

IFO pressure 1.7E-4 Torr on new not logged cold cathode gauge. P1 <7E-4 Torr

Valve configuration: vac.normal with anunulossess closed off.

TP3 was turned off with a failing drypump. It will be replaced tomorrow.

All time stamps are blank on the MEDM screens.

 

 

  13422   Thu Nov 9 15:33:08 2017 johannesUpdateCDSrevisiting Acromag
Quote:

We probably want to get a dedicated machine that will handle the EPICS channel serving for the Acromag system

http://www.supermicro.com/products/system/1U/5015/SYS-5015A-H.cfm?typ=H

This is the machine that Larry suggested when I asked him for his opinion on a low workload rack-mount unit. It only has an atom processor, but I don't think it needs anything particularly powerful under the hood. He said that we will likely be able to let us borrow one of his for a couple days to see if it's up to the task. The dual ethernet is a nice touch, maybe we can keep the communication between the server and the DAQ units on their separate local network.

ELOG V3.1.3-